[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Oct 09, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Oct 9 20:35:46 UTC 2017


HSU Shift Log for Mon Oct 09 2017 16:35:39 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: Model colors seem fine now. Colors by Intensity option was  
probably checked earlier.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: UW-CIMSS ADT values not available due to TS Ophelia  
not being posted on their web page.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171009/30c8b110/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list